OpenStudy (barreraa):

Evaluate the function at x = –3. (Points : 1) y = –3 y = 4 y = 3 y = 0

OpenStudy (barreraa):

@iGreen.

OpenStudy (yolo_king):

What do you think is the answer?

OpenStudy (barreraa):

C

OpenStudy (barreraa):

my best guess

OpenStudy (yolo_king):

C is not the answer.

OpenStudy (barreraa):

how about A

OpenStudy (anonymous):

the answer is 4 @barreraA

OpenStudy (barreraa):

is it 4

OpenStudy (barreraa):

so its B then

OpenStudy (yolo_king):

\[f(-3)=\sqrt{25-(-3)^2} =\sqrt{25-(9)}=\sqrt{25-9}=\sqrt{16}=4\]

OpenStudy (yolo_king):

B is the answer.